# Matching Non-Identical Objects

### v3

@@ 4 Proposed Method
In this section, we present a simple plug-and-play weighting module that can turn various sparse matchers to work for non-identical object matching. Most sparse matchers, such as [49, 27, 4, 38], consist of a pair of a keypoint detector and a feature matcher, and our module process the keypoints before feeding them to the feature matcher. Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, The keypoint detector outputs collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$. The proposed module then computes weighted descriptors $\widetilde{\bm{d}}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}$ for each $\bm{d}_{i}^{A},\bm{d}_{i}^{B}$, respectively. Lastly, the feature matcher performs matching using $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
4.1 Weighting module
The proposed weighting module uses an object detector and visual explanation model. This study used YOLOv7 [62] and Grad-CAM [51]. Let $\bm{x}\in{\mathcal{X}}$ be an image, and let $D=\{(\bm{p}_{i},\bm{d}_{i})\}_{i=1}^{n}$ be the collection of $n$ keypoints of $\bm{x}$.
First, the object detector is applied to $\bm{x}$ and detects the objects with their classes. For each of the detected objects (say, the $k$-th object), the visual explanation model gives pixel-wise class-activation scores or a heatmap $H_{k}$. Let $H_{k,\bm{p}_{i}}\in[0,1]$ be the heatmap score of $H_{k}$ at $\bm{p}_{i}$. Then, the heatmap scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$. Finally, the weighted descriptor $\tilde{\bm{d}}_{i}$ is computed by
@@ 4.1 Weighting module
We empirically found that the addition of base value 1 and the normalization by the maximum value is important. For example, the simplest approach $\widetilde{\bm{d}}_{i}=H_{\bm{p}_{i}}\bm{d}_{i}\in[0,1]^{d}$ ends up with a scarcity of matching because of the heatmap scores on the keypoints on background tend to be too low, i.e., $H_{\bm{p}_{i}}\approx 0$. Although such keypoints will not likely appear in the matching results, they seem to help the keypoints on objects to be matched. If one instead uses $\widetilde{\bm{d}}_{i}=(1+H_{\bm{p}_{i}})\bm{d}_{i}\in[1,2]^{d}$, this also affects negatively. We consider that this is because of the softmax operations used in the subsequent feature matching. We discuss the consequence of the weighting of descriptors next.
@@ 4.2 Effect of weights in feature matcher
where $I\in\{A,B\}$. Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as [^1] To be precise, this score should be considered as the pre-activation score, and the softmax operation will be subsequently applied. The terminology of this paper follows the related studies, e.g., [27].
@@
where $\mathcal{R}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{d\times d}$ denotes the rotary-encoding [53]. Therefore, our weighting by heatmap scores leads to the weighting of attention scores. The self-attention scores reflect the internal relationship of objects and background, thereby extracting better feature representations in the subsequent steps of the feature matcher. Roughly speaking, nothing that $\alpha_{i}^{I}\in[0.5,1.0]$ for all $i$ and $I\in\{A,B\}$, we have $\alpha_{i}^{I}\alpha_{j}^{I}\approx 1$ between the keypoints on objects, $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.5$ between one on object and the other on background, and $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0$ for between those on the background. This is reasonable because we want nice feature representations for object keypoints for nice matching. Nevertheless, such nice representations should be obtained by taking into account the background conditions, e.g., to compensate for the overall lighting conditions. If one uses $\alpha_{i}^{I}=H_{\bm{p}_{i}}$ as discussed earlier, i.e., $H_{\bm{p}_{i}}\approx 0$ for background keypoints, the feature representations cannot take into account the background.
@@
As discussed earlier, the value of $\alpha_{i}^{A}\alpha_{j}^{B}$ is generally large between the keypoints on objects. Clearly, this encourages the matching between objects.
4.3 Class Similarity Adapter
Figure 4: Class Similarity Adapter converts class labels from an object detector into embeddings using the CLIP [42] text encoder and compares them with cosine similarity to identify semantically similar object pairs. The GradCAM [51] heatmap of object detector YOLOv7 [62] generated for the chosen object pair is used for our weighting method, enhancing matching with multiple objects. Here, we used the keypoint detector SuperPoint [10] and the feature matcher LightGlue [27].
We propose the Class Similarity Adapter to address scenarios where multiple objects are present in an image. Our weighting module uses all heatmaps of objects detected by the object detector for weighting. However, this can needlessly emphasize irrelevant objects, leading to scattered matching from one to multiple objects. As shown in  Fig. 4 (bottom right), relying solely on visual information leads to matching based on similarities in texture, color, and shape, where similar black features such as the nose of the dog and the ear of the cat, are incorrectly matched, causing correspondences to scatter across multiple animals.
To resolve such confusion in the visual space, we again leverage semantic information given by the object detector. Particularly, we map the class labels of detected objects to text embedding space and determine the object pair to be matched based on the cosine similarity between embedding vectors. Suppose that an image $\bm{x}_{A}$ contains $n_{A}$ objects, and let $Y_{A}=\{y_{i}^{A}\}_{i=1}^{n_{A}}$ be their class labels. We define $Y_{B}=\{y_{i}^{B}\}_{i=1}^{n_{B}}$ similarly for the other image $\bm{x}_{B}$. Let $\mathcal{T}$ be a pretrained text encoder that maps a label $y$ to an embedding vector $\mathcal{T}(y)$. Particularly, we used the CLIP text encoder [42]. The Class Similarity Adapter determines the object pair $(i,j)$ to be matched based on the cosine similarity between the embedding vectors as follows.
@@ 4.3 Class Similarity Adapter
where $\|\,\cdot\,\|$ denotes the $L_{2}$ norm of vector. Then, the weighting module only uses the Grad-CAM heatmaps of the $i$-th object in $\bm{x}_{A}$ and the $j$-th object in $\bm{x}_{B}$. Based on this result, the weighting module uses the corresponding heatmaps for weighting the descriptors. Figure 4 shows this pipeline.
@@ 5 Experiments
We demonstrate the effectiveness of our method in non-identical object matching tasks, presenting qualitative results on various pairs of non-identical (but similar) objects, followed by quantitative evaluation on classical image matching tasks but under common corruptions [20].
Setup.
In the sparse matching framework, the image matching task is performed by a keypoint detector and a feature matcher. For the former, we adopted pre-trained SuperPoint [10], extracting 2,048 keypoints. For the latter, the extracted keypoints are matched using either pretrained SGMNet [4], GlueStick [38], and the state-of-the-art method, LightGlue [27], with and without the proposed weighting module. We additionally tested a detector-free dense matcher, LoFTR [56] and Efficient LoFTR [64]. We downloaded the pretrained models from their official repositories (details provided in the supplementary material). All experiments are conducted on a computer with an NVIDIA GeForce RTX 3090 GPU and 24GB memory.
@@ 5.1 Matching Non-Identical Objects
In this experiment, we address a matching of non-identical objects. It is worth noting that this task can only be qualitatively evaluated because there is no ground truth matching between non-identical objects.
Datasets.
@@
We collected various images with objects from the COCO [26] and ImageNet [9] datasets. To prepare the drawing version of several images, we used the DreamStudio [^2] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [47].
Figure 5: Typical categories of non-identical object matching. (a) A combination of SuperPoint [10] and LightGlue [27] only finds a small number of correspondence, and many of them are incorrect. (b) The proposed method significantly improves the matching.
Robust matching across non-identical objects.
@@
We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$). For all the cases, we have perspective difference (i.e., $P_{A}\neq P_{B}$). Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in Sec. A in the supplementary material. Figure 5 shows the matching results with and without the proposed method in the four cases. As can be seen in Fig. 5(a), a combination of SuperPoint and LightGlue cannot distinguish between objects and backgrounds or correspond to the same structure of animals. In contrast, Fig. 5(b) shows that the introduction of the proposed method significantly improves the results, reducing unreasonable matchings (e.g., human and racket, animal eye and mouth) and increasing reasonable part-to-part matchings. Particularly, for the cases (i) increased the number of matches between rackets and decreased the mismatches between people, (ii) corresponded to the same parts of animals, (iii) matched front wheels with domain gaps, and (iv) significantly increased the number of matches despite changes in background and object angles. Notably, structural features, such as eyes and noses, are successfully matched between different classes. We consider that the heatmaps of these animals gave similar importance to the eyes or mouth, thus increasing the number of matches between the same parts of the body. In addition, the results for the same class and appearance showed a decrease in mismatches with the background, which means that the background and foreground had relatively lower similarity.
Different matchers.
@@
Figure 6 presents the results with other matchers, showing that our method boosts all of them to match corresponding parts between different animals.
@@
Table 1: The pose accuracy (AUC) at the maximum angular error of $20\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions [20]. The left side shows both images corrupted with the same type of corruption, while the right side shows only one of the input images corrupted and the other as a clean image. Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC. The dense matchers, LoFTR [56] and Efficient LoFTR (ELoFTR) [64], are shown for reference. Our method improves the robustness of sparse matchers with a slight reduction in clean accuracy and runtime.
@@
LG LG+Ours GS GS+Ours LoFTR ELoFTR LG LG+Ours GS GS+Ours LoFTR ELoFTR

### v4

@@ 4 Proposed Method
We propose a plug-and-play module, Semantic Enhancement Weighting (SEW), that can be inserted into various sparse matchers, such as SuperGlue [69], LightGlue [37], and more [7, 54]. Sparse matchers consist of a pair of a keypoint detector and a feature matcher. Our module processes the keypoints before feeding them to the feature matcher. Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, The keypoint detector outputs collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$. The proposed module then computes weighted descriptors $\widetilde{\bm{d}}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}$ for each $\bm{d}_{i}^{A},\bm{d}_{i}^{B}$, respectively. Lastly, the feature matcher performs matching using $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
As will be shown in Sec. 5, this simple pipeline works surprisingly well and adapts pretrained sparse matching models without any training, addressing the technical challenge in non-identical object matching discussed above.
4.1 Semantic Enhancement Weighting
The proposed weighting module uses an object detector and a visual explanation model. This study used YOLOv7 [87] and Grad-CAM [73]. Let $\bm{x}\in{\mathcal{X}}$ be an image, and let $D=\{(\bm{p}_{i},\bm{d}_{i})\}_{i=1}^{n}$ be the collection of $n$ keypoints of $\bm{x}$.
First, the object detector is applied to $\bm{x}$ and detects the objects with their classes. For each of the detected objects (say, the $k$-th object), the visual explanation model gives pixel-wise class-activation scores or a heatmap $H_{k}$. Let $H_{k,\bm{p}_{i}}\in[0,1]$ be the heatmap score of $H_{k}$ at $\bm{p}_{i}$. Then, the scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$. [^1] The max operation outperformed other candidates including mean, sigmoid, and softmax operations in our preliminary experiments. Finally, the weighted descriptor $\tilde{\bm{d}}_{i}$ is computed by
@@ 4.1 Semantic Enhancement Weighting
We empirically found that the addition of base value 1 and the normalization by the maximum value is important. For example, the simplest approach $\widetilde{\bm{d}}_{i}=H_{\bm{p}_{i}}\bm{d}_{i}\in[0,1]^{d}$ ends up with a scarcity of matching because of the heatmap scores on the keypoints on background tend to be too low, i.e., $H_{\bm{p}_{i}}\approx 0$. Although such keypoints will not likely appear in the matching results, they seem to help the keypoints on objects to be matched. If one instead uses $\widetilde{\bm{d}}_{i}=(1+H_{\bm{p}_{i}})\bm{d}_{i}\in[1,2]^{d}$, this also affects negatively. We consider that this is because of the softmax operations used in the subsequent feature matching. We next discuss how the weighting of descriptors impacts the matching.
@@ 4.2 Effect of weights in feature matcher
where $I\in\{A,B\}$. Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as
@@
where $\mathcal{R}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{d\times d}$ denotes the rotary-encoding [76]. Therefore, our weighting by heatmap scores leads to the weighting of attention scores. The self-attention scores reflect the internal relationship of objects and background, thereby extracting better feature representations in the subsequent steps of the feature matcher. Roughly speaking, noting that $\alpha_{i}^{I}\in[0.5,1.0]$ for all $i$ and $I\in\{A,B\}$, we have $\alpha_{i}^{I}\alpha_{j}^{I}\approx 1$ between the keypoints on objects, $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.5$ between one on object and the other on background, and $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.25$ for between those on the background. This is reasonable because we want nice feature representations for object keypoints for nice matching. Nevertheless, such nice representations should be obtained by taking into account the background conditions, e.g., to compensate for the overall lighting conditions. If one uses $\alpha_{i}^{I}=H_{\bm{p}_{i}}$ as discussed earlier, i.e., $H_{\bm{p}_{i}}\approx 0$ for background keypoints, the feature representations cannot take into account the background.
@@
As discussed earlier, $\alpha_{i}^{A}\alpha_{j}^{B}$ is large between keypoints of the objects, encouraging matching between objects.
4.3 Non-visual Object Pairing
We next propose Non-visual Object Pairing for the case where images contain multiple objects. In such a case, it is essential to determine which object in one image should be matched with another object in the other image, and the weighting module should exploit the heatmaps of them and exclude those of irrelevant objects. Otherwise, the local spurious similarity of objects causes scattered and inconsistent matching; see  Fig. 5 (bottom right), where a nose of a dog and an ear of a cat match.
A pair of similar objects for matching can be determined using visual or non-visual features. For the former, we can use the feature vectors extracted from an intermediate layer of the object detector. This can be done simultaneously during the heatmap computation process. For the latter, text embedding (e.g., by the CLIP text encoder) of object labels may be a reasonable choice. The text embedding vectors can be pre-computed for all the classes if one pursues fast matching. The visual and non-visual features are complementary, and the choice should depend on the use case. For example, given (lesser panda, raccoon, giant panda), the visually similar pair is (lesser panda, raccoon), while (lesser panda, giant panda) is more conceptually similar.
Figure 5: Non-visual Object Pairing converts class labels from an object detector into embeddings using the CLIP [59] text encoder and compares them with cosine similarity to identify semantically similar object pairs. The GradCAM [73] heatmap of object detector YOLOv7 [87] generated for the chosen object pair is used for our weighting method, enhancing matching with multiple objects.
We here provide concrete procedures in the Non-visual Object Pairing based on non-visual features. Suppose that an image $\bm{x}_{A}$ contains $n_{A}$ objects, and let $Y_{A}=\{y_{i}^{A}\}_{i=1}^{n_{A}}$ be their class labels. We define $Y_{B}=\{y_{i}^{B}\}_{i=1}^{n_{B}}$ similarly for the other image $\bm{x}_{B}$. Let $\mathcal{T}$ be a pretrained text encoder that maps a label $y$ to an embedding vector $\mathcal{T}(y)$. Particularly, we used the CLIP text encoder [59]. The adapter determines the object pair $(i,j)$ to be matched based on the cosine similarity between the embedding vectors as follows.
@@ 4.3 Non-visual Object Pairing
[…]
@@ 5 Experiments
In this section, we demonstrate the effectiveness of the proposed method in non-identical object matching tasks. Section 5.1 presents various qualitative comparisons from state-of-the-art sparse feature matching methods. Section 5.2 then introduces Triangular Matching Consistency (TMC), a new annotation-free metric that allows us to quantitatively evaluate non-identical object matching itself. Section 5.3 presents quantitative evaluations. In the latter experiment, as there are no available datasets providing fine-grained matching between non-identical objects, we examine the matching results between identical objects under 15 types of corruptions, which measures the robustness to domain shifts. We emphasize that while semantic correspondence is a conceptually similar task to ours, it does not target fine-grained matching, and thus, our experiment setup (i.e., datasets, baseline methods, and metrics) is based on feature matching literature.
Figure 6: Typical categories of non-identical object matching. (a) A combination of SuperPoint [16] and LightGlue [37] only finds a small number of correspondences, and many of them are incorrect. (b) Our method SEW significantly improves the matching.
Setup. The sparse matching pipeline involves a keypoint detector and a feature matcher. For the former, we adopted pre-trained SuperPoint [16], extracting 2,048 keypoints. For the latter, the extracted keypoints are matched using either pretrained SGMNet [7], GlueStick [54], and the state-of-the-art method, LightGlue [37], with and without the proposed method. We additionally tested a detector-free dense matcher, LoFTR [79] and Efficient LoFTR [89]. We downloaded the pretrained models from their official repositories (details provided in the supplementary material). All experiments used NVIDIA GeForce RTX 3090 GPU and 24GB memory.
@@ 5.1 Matching Non-Identical Objects
Here, we perform the matching of non-identical objects and evaluate them qualitatively.
@@
Datasets. We collected various images with objects from the COCO [36] and ImageNet [15] datasets. To prepare the drawing version of several images, we used the DreamStudio [^2] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [66].
@@
Robust matching across non-identical objects. We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$). For all the cases, we have a perspective difference (i.e., $P_{A}\neq P_{B}$). Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in Sec. A in the supplementary material. Figure 6 shows the matching results with and without the proposed method in the four cases. As shown in Fig. 6 (a), SuperPoint and LightGlue alone fail to distinguish objects from the background or match corresponding animal structures. In contrast, Fig. 6 (b) shows that the proposed method significantly reduces unreasonable matches (e.g., human and racket, animal eye and mouth) while increasing correct part-to-part matchings. Particularly, it (i) increased racket matches while reducing mismatches between people, (ii) aligned corresponding animal parts, (iii) matched front wheels despite domain gaps, and (iv) significantly increased matches despite background and angle variations. Notably, structural features, such as eyes and noses, are successfully matched between different classes. We consider that the heatmaps of these animals gave similar importance to the eyes or mouth, thus increasing the number of matches between the same parts of the body. In addition, the results for the same class and appearance showed a decrease in mismatches with the background, which means that the background and foreground had relatively lower similarity.
@@
Different matchers. Figure 7 presents the results with other matchers, showing that our method boosts all of them to match corresponding parts between different animals.
Figure 7: The results with different matchers [7, 54, 37]. SuperPoint [16] is used for the keypoint detection. The introduction of SEW increases the matching of eyes and noses between the cheetah and husky and decreases mismatching in the background.
5.2 Triangular Matching Consistency
Direct quantitative evaluation of non-identical object matching is difficult, as fine-grained ground-truth correspondences between non-identical objects are inherently ambiguous. We therefore introduce Triangular Matching Consistency (TMC), an annotation-free metric. TMC measures the agreement between direct matching $(A,C)$ and one-hop matching via $(A,B)$ and $(B,C)$, where $A,B,C$ are semantically related images. By selecting $(A,C)$ to be highly similar, their direct matches provide a reliable reference, and a robust our task method should approximate them even through the one-hop path. TMC reports error magnitude (RMSE), accuracy (PCK), and coverage (Recall). See Sec. E and Fig. C for the details.
Dataset. We build an image set of 1,500 triplets for TMC. We use animal images from the Wild category of the Animal Faces dataset [11]. This dataset contains diverse animal face images, including dogs, cats, lions, tigers, and more. For each image $A_{i}$, we generate a new view $C_{i}$ of the same object by applying a 0.9× pinch-out geometric transformation. Then, another species image $A_{j}$ ($j\neq i$) is used as $B$, forming $(A,B,C)=(A_{i},A_{j},C_{i})$. Here, $(A,C)$ are identical-instance pairs with viewpoint change, while $B$ provides a semantically similar but different object. This design allows us to use $(A,C)$ as a reliable reference while introducing non-identical variation through $B$, which is essential for evaluating our new task.
Results. Table 1 summarizes the evaluation. RMSE measures geometric error, and our module reduces it to about half of each baseline. It even outperforms dense detector-free methods. PCK measures the proportion of matches within a tolerance. It improves at all thresholds, with the largest gains at broader ones and clear benefits even under strict settings. This shows that our method enhances consistent matching between corresponding object parts. Recall measures how well one-hop correspondences cover the $A{\leftrightarrow}C$ baseline, and it also increases with our module. Across all three metrics, it indicates that our method enhances the matching consistency of non-identical objects.
Table 1: Triangular Matching Consistency (TMC) results for the wild category in the Animal Faces dataset [11]. Sparse matchers (LightGlue (LG), GlueStick (GS)) with SuperPoint; dense matchers (LoFTR, ELoFTR) are shown for reference. Metrics include RMSE (lower is better), PCK, and Recall at distance thresholds of 0.01, 0.05, and 0.10 of the image size. Our method SEW shows large improvements across all metrics. See Sec. F for details.
[…]
0.01 0.05 0.10 0.01 0.05 0.10
Sparse Matcher (with SuperPoint) LG 0.14 54.0 68.9 68.2 70.2 76.9 81.4
LG+SEW 0.07 73.8 87.8 92.6 73.8 87.7 92.9
GS 0.16 49.0 62.1 62.3 63.8 71.3 73.0
GS+SEW 0.08 68.8 81.0 86.6 67.3 82.1 86.3
Dense Matcher (detector-free) LoFTR 0.19 52.7 61.9 66.8 57.3 62.1 65.7
ELoFTR 0.14 64.5 73.7 76.7 65.9 73.3 76.9
@@ 5.2 Triangular Matching Consistency
Table 2: The pose accuracy (AUC) at the maximum angular error of $20\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [79] under common corruptions [27]. The left side shows both images corrupted with the same type of corruption, while the right side shows only one of the input images corrupted and the other as a clean image. Our method SEW largely improves LightGlue (LG) [37] and GlueStick (GS) [54] for most categories and the average AUC. The dense matchers, LoFTR [79] and Efficient LoFTR (ELoFTR) [89], are shown for reference. Our method improves the robustness of sparse matchers with a slight reduction in clean accuracy and runtime.
@@
LG LG+SEW GS GS+SEW LoFTR ELoFTR LG LG+SEW GS GS+SEW LoFTR ELoFTR